a person is moving in same direction as muon measures lifetime of muon to be 1 micro second if speed of muon and person with respect to lab are c/2 and c/3 .calculate proper life time of muon and distance moved by muon with respect to lab.
1
Expert's answer
2020-09-08T10:51:37-0400

solution:-

given data

speed of muon with respect to lab(v)="\frac{c}{2}"

speed of person with respect to lab(v')="\frac{c}{3}"

relative velocity can be written as


"v_{rel}=\frac{v-v'}{1-\frac{vv'}{c^2}}"


"v_{rel}=\frac{\frac{c}{2}-\frac{c}{3}}{1-\frac{c^2}{6c^2}}=\frac{c}{5}"

proper life time for muon can be written


"\Delta t'=\Delta t \sqrt{1-\frac{v_{rel}^2}{c^2}}"

by putting the value


"\Delta t=1\sqrt{1-\frac{c^2}{25c^2}}=0.9797\mu s"


distance moved by muon with respect to lab is


"d=v_{rel} \Delta t'=\frac{c}{5}\times(0.9797)\times10^{-6}"


"=\frac{3\times10^8}{5}\times(0.9797)\times10^{-6}"


"d=58.78m"


therefore distance moved by muon is 58.78m.